Solve the logarithmic equation 23x-2 =32x-1

1
Expert's answer
2021-11-29T13:22:31-0500
"2^{3x-2} =3^{2x-1}"

"\ln (2^{3x-2})=\ln (3^{2x-1})"

"(3x-2)\ln (2)=(2x-1)\ln (3)"

"x(3\ln (2)-2\ln(3))=2\ln (2)-\ln(3)"

"x=\dfrac{\ln(8/9)}{\ln(4/3)}"